# Efficient Causal Graph Discovery Using Large Language Models

### v3

@@ 4.3 Results
Tables 1, 2 and 3 show the results of our experiments on the Asia causal graph, the Child causal graph, and the Neuropathic Pain causal graph respectively. We bold the best result in the F Score column and NHD Ratio column since those are the main metrics we are looking at for comparison.
@@
On this much larger graph, we see that even the methods which are able to run fail catastrophically, barely outperforming the baseline which predicts the same number of edges but all wrong (as seen by the NHD Ratios which are all above 0.9). In contrast, the proposed method achieves reasonable performance with an F Score of 0.351 and a NHD Ratio of 0.643, thus emerging as the only method which is able to handle larger graphs such as this one.
@@ 7 Future Work
Another avenue to explore is using advanced prompting strategies like Tree of Thoughts (Yao et al., 2023), which are known to significantly improve performance in complex multi-step reasoning tasks like causal graph discovery.
@@
8 Impact Statement
Causal graph discovery has many potential applications in medicine (understanding Alzheimer’s pathophysiology), biology (constructing gene regulatory networks), and psychology (understanding the causal relationships between psychiatric disorders) (Zanga & Stella, 2023) which, if successful, could have far-reaching positive impacts on society. We hope that our work takes a step towards achieving those positive impacts by advancing the field of causal graph discovery.

### v4

@@ 4.3 Results
Tables 1 shows the results of our experiments on the Asia causal graph, the Child causal graph, and the Neuropathic Pain causal graph respectively. We bold the best result in the F Score column and NHD Ratio column since those are the main metrics we are looking at for comparison.
@@
On this much larger graph, we see that even the methods that are able to run fail catastrophically, barely outperforming the baseline which predicts the same number of edges but all wrong (as seen by the NHD Ratios which are all above 0.9). In contrast, the proposed method achieves reasonable performance with an F Score of 0.351 and a NHD Ratio of 0.643, thus emerging as the only method which is able to handle larger graphs such as this one.
@@ 7 Future Work
Another avenue to explore is using advanced prompting strategies like Tree of Thoughts Yao et al. (2023), which are known to significantly improve performance in complex multi-step reasoning tasks like causal graph discovery.
@@
8 Reproducibility Statement
We provide code to reproduce all experiments in this anonymized repository
9 Acknowledgements
We are grateful to Dhanya Sridhar, Philippe Brouillard, and Marco Jiralerspong for their helpful ideas and suggestions.